OpenStudy (anonymous):

Tammy, Chang, and Deandre have a total of $123 in their wallets. Deandre has $9 more than Tammy. Chang has 4 times what Deandre has. How much do they have in their wallets?

OpenStudy (imstuck):

you need to base all the amounts in their wallets on a single variable, so we will use x. It appears to me that poor Tammy has the least amount in her wallet, because it tells me that Deandre has 9 more than Tammy, and Chang has even more than Deandre does. So we will say that what Tammy has in her wallet is x. Then Deandre has 9 more, so that's x + 9. Chang has 4 times what Deandre has so that's 4(x + 9). All the money collectively equals $123. Here's the set up, then:

OpenStudy (imstuck):

T - x D - x + 9 C - 4(x + 9) x + (x + 9) + [4(x + 9)] =

OpenStudy (imstuck):

That simplifies to: x + x + 9 + 4x + 36 = 123 6x + 45 = 123 6x = 78 x = 13 That means that Tammy has $13 in her wallet; Deandre has 13 + 9 = $22; Chang has 4(22) = $88. Add them all together to make sure they equal $123: 13+ 22 + 88 = $123

OpenStudy (anonymous):

thank you for helping me understand

OpenStudy (imstuck):

Did you understand ok? If you have questions on any of this, I will be happy to explain.

OpenStudy (anonymous):

Yes! And Okay